OpenStudy (anonymous):

i cant seem to figure out this equation, 4(b-3)-b=7b+10

OpenStudy (anonymous):

have you tried solving it?

OpenStudy (shinalcantara):

you try to multiply 4 to the quantity inside the parenthesis.

OpenStudy (anonymous):

yes i have tries to solve it, i will try multiplying by 4

OpenStudy (shinalcantara):

what you've got?

OpenStudy (anonymous):

i still cant figure it out

OpenStudy (anonymous):

show us your solution so that we can check it

OpenStudy (anonymous):

i need to find what b is, thats what i cant figure out

OpenStudy (anonymous):

4(b-3)-b=7b+10 4b-12-b=7b+10 3b-12=7b+10 3b-7b=10+12 -4b=22 b=22/-4 b=-11/2

OpenStudy (anonymous):

i hope you understand what @secretslocked did. But remember that we are just here to GUIDE you, not to do the homework for you.

OpenStudy (anonymous):

yes i understand, but thats calculus, im in algebra 1

OpenStudy (anonymous):

i dont see any calculus here? O.o that's an algebraic equation, which you can solve algebraically. therefore i can say that, that is part of algebra:)

OpenStudy (anonymous):

sorry, its over my head

OpenStudy (anonymous):

it's okay. I'm just saying ^_^

OpenStudy (anonymous):

i need simple help lol, not complicated help lol

OpenStudy (anonymous):

:)